# Three-dimensional Morphological Reconstruction of Millimeter-Scale Soft Continuum Robots based on Dual Stereo Vision

### v1

@@ II-A Capturing Raw 3D Point Clouds
To capture the raw 3D point clouds for the NTCR, as shown in Fig. 2, we built a dual-stereo vision system by utilizing two Intel RealSense D405 cameras. These two cameras are arranged at a calibrated baseline distance to capture the opposite perspectives of the object (i.e., the small continuum robot). Although the system can become more explicit by using multiple cameras [16], we opt for the minimal required visions to reduce the data redundancy. These cameras generate depth maps, which provide critical spatial information about the object’s shape and rough appearance. The depth $Z$ of a point $P$ can be derived from the disparity $d$ between the corresponding points in the left and right camera images, i.e.,
@@ II-B Conversion and Filtering
The conversion and filtering stage is pivotal for refining raw data obtained from 3D scanning to produce accurate 3D models. Depth maps captured by the dual stereo vision system are first converted into point cloud data [20]. The transformation from pixel coordinates $(u,v)$ to camera coordinates $(X_{c},Y_{c},Z_{c})$ is given by the intrinsic camera matrix $K$:
@@ III-A Experiment Setup
The experimental setup for 3D scanning and reconstruction of millimeter-scale NTCR involves a custom-designed sliding track system equipped with two Intel RealSense D405 depth cameras. The cameras are mounted on adjustable slides that can move along the track, facilitating the calibration and optimization of their relative positions. A 3D-printed rack is placed in between the cameras that allows the NTCR to maintain a fixed position at the robot base, and this base frame can be referred to as the reference frame. The base features a vertical post with small holes for anchoring the robot, ensuring its stability and preventing movement during scanning. This arrangement is crucial for maintaining consistent reference coordinates throughout the scanning process, particularly when the robot’s posture is altered by manipulating an internal control wire. As such, a point cloud data collection platform is established. The setup is shown in Figure 2. This setup ensures that the reference coordinates of the robot remain consistent, even when its configuration is altered by manipulating an internal control wire.

### v3

# Three-dimensional Morphological Reconstruction of Millimeter-Scale Soft Continuum Robots based on Dual-Stereo-Vision

@@ II-A Capturing Raw 3D Point Clouds
To capture the raw 3D point clouds for the NTCR, as shown in Fig. 2, we built a dual-stereo-vision system by utilizing two Intel RealSense D405 cameras. These two cameras are arranged at a calibrated baseline distance to capture the opposite perspectives of the object (i.e., the small continuum robot). Although the system can become more explicit by using multiple cameras [16], we opt for the minimal required visions to reduce the data redundancy. These cameras generate depth maps, which provide critical spatial information about the object’s shape and rough appearance. The depth $Z$ of a point $P$ can be derived from the disparity $d$ between the corresponding points in the left and right camera images, i.e.,
@@ II-B Conversion and Filtering
The conversion and filtering stage is pivotal for refining raw data obtained from 3D scanning to produce accurate 3D models. Depth maps captured by the dual-stereo-vision system are first converted into point cloud data [20]. The transformation from pixel coordinates $(u,v)$ to camera coordinates $(X_{c},Y_{c},Z_{c})$ is given by the intrinsic camera matrix $K$:
@@ III-A Experiment Setup
The experimental setup for 3D scanning and reconstruction of millimeter-scale NTCR involves a custom-designed sliding track system equipped with two Intel RealSense D405 depth camerasThe cameras are mounted on adjustable slides that can move along the track, facilitating the calibration and optimization of their relative positions. A 3D-printed rack is placed in between the cameras that allows the NTCR to maintain a fixed position at the robot base, and this base frame can be referred to as the reference frame. The base features a vertical post with small holes for anchoring the robot, ensuring its stability and preventing movement during scanning. This arrangement is crucial for maintaining consistent reference coordinates throughout the scanning process, particularly when the robot’s posture is altered by manipulating an internal control wire. As such, a point cloud data collection platform is established. The setup is shown in Figure 2. This setup ensures that the reference coordinates of the robot remain consistent, even when its configuration is altered by manipulating an internal control wire.